A stdy involves 608 randomly selected deaths, with 26 of them caused by accidents. Construct a 98% confidence interval for the true percentage of all deaths that are caused by accidents.
i now the answer is 2.36%
Answer by stanbon(57290)

A stdy involves 608 randomly selected deaths, with 26 of them caused by accidents. Construct a 98% confidence interval for the true percentage of all deaths that are caused by accidents.
I know the answer is 2.36%
--------------------------------------
sample proportion = 26/608 = 0.043
standard error = E = 2.3263*sqrt[(0.043)(0.957)/608] = 0.019..
-----------------------------------
98% CI: 0.043-0.019 < p < 0.043+0.019
---
98% CI: 0.024 < p < 0.062
========================================
Cheers,
Stan H.